if 2y x 4x + 40 = 180 then what is the value of 2x + y ... thanks
Answer by drk(1908)

We have the original equation as
(i) 2y x 4x + 40 = 180
subtract 40 to get
8xy = 140
(ii) xy = 17.5 = 35/2
(iii) 2x = 35y
(iv) 4x = 70y
substituting back into (i) and subtracting 40, we get
(v) 2y*70y = 140
(vi) 140y^2 = 140.
Y = +-1.
If y = 1, then x = 17.5
If y = -1, then x = -17.5
They both work in the original equation.